1.Find the centroid of the region bounded by the curve 𝑦 = π‘₯3 and 𝑦 = 4π‘₯ in the first quadrant.Β 


2. Find the centroid of the solid of revolution generated by revolving the region bounded by 𝑦 = 4π‘₯ βˆ’ π‘₯2 and the x-axis about the y-axis.


1
Expert's answer
2021-05-24T08:01:02-0400

1

"Mx=\frac{1}{2}\int_a^b(f(x)^2-g(x)^2)dx=\frac{1}{2}\int_0^1[(4x)^2-(x^3)]"

"=2.59523"

"My=\int_a^bx[f(x)-g(x)]dx=\int_0^1[x[(4x)^2-(x^3)dx]"

"=\:3.85714"

"M=\int_a^b[f(x)-g(x)]dx=\int_0^1[4x-x^3]"

"=1.75"

"\bar{x},\bar{y}=(\frac{My}{M},\frac{Mx}{M})=(2.20408,1.48289)"

2

"y=4x-x^2"

"4x-x^2=x"

"x=0,3"

we are interested in the point in the first quadrant, the transition point is

"x_t,y_t=3,3"

"\bar{y}=\int\frac{ydv}{v}" where v = volume

"V=\int dV=\int \pi x^2dy"

"V=\int_0^{y_t}\pi^2dy+\int_{y_t}^4 \pi(\sqrt{4-y)}dy"

"V=\pi[(\frac{1}{3}y^3)^{y_t}_0+(4y-\frac{1}{2}y^2)]^4_{y_t}"

"V=\frac{\pi}{12}(121-17\sqrt{17)}"

"\int ydV=\int y(\pi x^2 dy)"

"\int ydV=\int_0^{y_t}\pi^3dy+\int_{y_t}^4\pi y(4-y)dy"

"\int ydV=\pi [(\frac{1}{4}y^4)_0^{y_t}+(2y^2-\frac{1}{3}y^3)_{y_t}^4]"

"\int ydV=\frac{x}{24}(135+17\sqrt{17)}"

finally we find the the centroid:

"\bar{y}=\frac{\frac{x}{24}(135+17\sqrt{17)}}{\frac{x}{12}(121-17\sqrt{17)}}"

"\bar{y}=\frac{83+17\sqrt{17}}{76}\approx 2.014379"